OpenStudy (anonymous):

Which pair of angles has congruent values for the sin x° and the cos y°? (4 points) Select one: a. 40°; 140° b. 40°; 80° c. 40°; 50° d. 40°; 40°

OpenStudy (anonymous):

@TuringTest @timo86m @Opcode @phi @ganeshie8

OpenStudy (raden):

hint : cos(y) = sin(90 - y)

OpenStudy (anonymous):

what would y be

OpenStudy (raden):

take it one of in options, wht do you get ?

OpenStudy (anonymous):

is it B

OpenStudy (raden):

let's recheck again, if the answer is B b. 40°; 80° it mean, x = 40° and y = 80° is sin(40) = cos80 ? by using the identity above, cos(y) = sin(90 - y) put y = 80 to both sides, get cos(80) = sin(90 - 80) cos(80) = sin(10) obvious, sin 40 not equals sin(10). so it is not right answer please, try again

OpenStudy (anonymous):

Ok its C i get it now

OpenStudy (raden):

yes, you are right !

OpenStudy (anonymous):

i see what you did there

OpenStudy (anonymous):

can you fan me please

OpenStudy (raden):

actually, the asker be fan me ;) lol

OpenStudy (anonymous):

i did

OpenStudy (anonymous):

can you help with another